# Tianyan: Cloud services with quantum advantage

### v1

@@ VIII Data availability
Hantao Sun1∗, Xinfang Zhang1∗, Zhen Wang1, Zilong Zha1, Jie Chen1, Qiankun Wang1, Jiasheng Hu1, Hanyi Wang1, Jianjian Gao1, Junhe Wang1, Bihao Guo1, Chunwang Liu1, Yang Li1, Yu Fan1, Jian Ding1, Jiping Fu1, Yuanhang Hu1, Rui Yang1, Weizhi Tao1, Zhehui Wang2, Huijie Guan2, ShuXin Xie2, Liang Zhou2, Jiafei Wang2, LiangChao Sun2, Zhong Zhao2, Chang Wang2, WenHao Chu2, Shengbin Wang1, Tao Yuan1, Xixi Zhang1, Lei Miao1, Wei Liu1, Wenya Huang1, Mengtian Zhou1, Hanze Guo1, He Huang1, Yang Jin1, Yin Kan1, Zhaowei Li1, Zirun Luo1, Yuan Liu1, Lijun Sun1, Yue Tian1, Mian Tao1, Jiawang Wan1, Rui Xu1, Jiahao Xie1, Shulin Yu1, Yuting Zhang1, Shuyue Zheng1, Canxia Zhou1, Tianqi Zong1

### v2

@@ VIII Data availability
Hantao Sun1,2∗, Xinfang Zhang1,2∗, Zhen Wang1,2, Zilong Zha1,2, Jie Chen1,2, Qiankun Wang1,2, Jiasheng Hu1,2, Hanyi Wang1,2, Jianjian Gao1,2, Junhe Wang1,2, Bihao Guo1,2, Chunwang Liu1,2, Yang Li1,2, Yu Fan1,2, Jian Ding1,2, Jiping Fu1,2, Yuanhang Hu1,2, Rui Yang1,2, Weizhi Tao1,2, Zhehui Wang3, Huijie Guan3, ShuXin Xie3, Liang Zhou3, Jiafei Wang3, LiangChao Sun3, Zhong Zhao3, Chang Wang3, WenHao Chu3, Shengbin Wang1,2, Tao Yuan1,2, Xixi Zhang1,2, Lei Miao1,2, Wei Liu1,2, Wenya Huang1,2, Mengtian Zhou1,2, Hanze Guo1,2, He Huang1,2, Yang Jin1,2, Yin Kan1,2, Zhaowei Li1,2, Zirun Luo1,2, Yuan Liu1,2, Lijun Sun1,2, Yue Tian1,2, Mian Tao1,2, Jiawang Wan1,2, Rui Xu1,2, Jiahao Xie1,2, Shulin Yu1,2, Yuting Zhang1,2, Shuyue Zheng1,2, Canxia Zhou1,2, Tianqi Zong1,2
@@
2 Quantum Research Institute of China Telecom, Hefei 230031, China